13. Sulphur is a raw material for the manufacture of sulphuric acid in Contact process


a) Describe the steps involved in the manufacture of sulphurlc acid by the 'Contact'process. In each step, write a balanced chemical equation for the reaction that occurs.

b) State two large scale uses of sulphuric acid 

c) How is the manufacture of sulphuric acid likely to harm environment.



14. A compound contains 85.7% by mass of carbon and 14.3%by mass of hydrogen. Its molar mass is 56g/mol. (C= 12, H=1)a) a)Calculate its empirical formula. 

b) Determine its molecular formula.

c) Draw a possible structural formula of the compound.

d) Write a balanced chemical equation for its combustion in oxygen.


15. A sample of lead (II) iodide (an insoluble salt) was prepared by reacting2O cm3 of O.2 mol dm-3 lead (II) nitrate with 40 cm3 of 0.2 mo1 dm-3Potassium iodide

a) What type of reaction occurs between lead (II) nitrate and potassium iodide solutions?

b) Calculate the number of moles of lead (II) nitrate in 20 cm3 of the solution. 

c) Write a balanced equation for the reaction between lead (II) nitrate and potassium iodide solutions.

d) Calculate the mass in grams of lead (II) iodide formed.(Pb=207, I= 127)


 16. a) Using specific examples, distinguish between a strong acid and a weak acid. 

b) Describe a simple laboratory test you could use to distinguish between a strong acid and a weak acid. State the reagent and the observation.


17. In upper atmosphere, there is a layer of ozone surrounding the earth.

a) Explain the importance of this layer in terms of human health.

b) State the type of chemical substance that destroy the ozone layer.


18. Name the process by which the components of the following mixtures can be separated:

a) Water and ethanol

b) Pigments and leaves

c) Sand and water

d) Ammonium chloride and sodium chloride.   


19. Nicotine, an alkaloid in the nightshade family of plants that is mainly responsible for the addictive nature of cigarettes, contains 74.02% C, 8.710% H, and 17.27% N. If 40.57 g of nicotine contains 0.2500 mol nicotine, what is the molecular formula?  


1
Expert's answer
2021-02-12T04:56:18-0500

13. a) The Contact Process is used in the manufacture of sulfuric acid. ... Step 1: Make sulfur dioxide. Step 2: Convert sulfur dioxide into sulfur trioxide (the reversible reaction at the heart of the process) Step 3: Convert sulfur trioxide into concentrated sulfuric acid.

Step 1: S + O2 = SO2

Step 2: 2SO2 + O2 = 2SO3

Step 3: SO3 + H2O = H2SO4.

b) The major use of sulfuric acid is in the production of fertilizers, e.g., superphosphate of lime and ammonium sulfate. It is widely used in the manufacture of chemicals, e.g., in making hydrochloric acid, nitric acid, sulfate salts, synthetic detergents, dyes and pigments, explosives, and drugs.

c) Sulfuric acid will exist as particles or droplets in the air if released to the atmosphere. It dissolves when mixed with water. It has moderate acute (short-term) toxicity on aquatic life. Sulfuric acid is very corrosive and would badly burn any plants, birds or land animals exposed to it.

14. C2H4.

H-C=C-H

|H H|